# Realistic predictions for Gaia black hole discoveries: comparison of isolated binary and dynamical formation models

### v1

@@ Abstract
Astrometry from Gaia has enabled discovery of three dormant black holes (BHs) in au-scale binaries. Numerous models have been proposed to explain their formation, including several that have forecasted Gaia detections. However, previous works have used simplified detectability metrics that do not capture key elements of the Gaia astrometric orbit selection function. We apply a realistic forward-model of Gaia astrometric orbit catalogs to BH binary populations generated through (a) isolated binary evolution (IBE) and (b) dynamical formation in star clusters. For both formation channels, we analyze binary populations in a simulated Milky Way-like galaxy with a realistic metallicity-dependent star formation history and 3D dust map. We generate epoch astrometry for each binary from the Gaia scanning law and fit it with the cascade of astrometric models used in Gaia DR3. The IBE model of Chawla et al. (2022) predicts that no BH binaries should have been detected in DR3 and thus significantly underpredicts the formation rate of Gaia BHs. In contrast, the dynamical model of Di Carlo et al. (2024) overpredicts the number of BHs receiving DR3 orbital solutions by a factor of 8. The two models predict very different orbital period distributions, with the IBE model predicting only binaries that avoided common envelope evolution and have d to be detectable, and the dynamical formation model predicting a period distribution that is roughly log-uniform. Adopting the dynamical channel as a fiducial model and rescaling by a factor of 1/8 to match DR3, we predict that 30 BH binaries will be detected in Gaia DR4, representing of Milky Way BHs with luminous companions in au-scale orbits.
@@ 1 Introduction
The formation of Gaia BH binaries remains uncertain. Isolated binary evolution (IBE) models struggle to reproduce the systems’ wide orbits. Their current orbital separations are significantly smaller than the predicted maximum radii of their BH progenitors, suggesting that their luminous stars would have interacted with the progenitors when they were red supergiants. However, these binaries would also have to survive an episode of common-envelope evolution with a large donor-to-accretor mass ratio; even if the envelopes of the massive stars were successfully ejected, the final orbits are predicted to be far tighter than observed. Modified prescriptions of IBE (e.g., Kruckow et al., 2024; Gilkis & Mazeh, 2024; Kotko et al., 2024), formation in a hierarchical triple (e.g., Nagarajan et al., 2024; Generozov & Perets, 2024; Li et al., 2024) and dynamical assembly in a star cluster (e.g., Rastello et al., 2023; Tanikawa et al., 2024; Di Carlo et al., 2024; Fantoccoli et al., 2024) have all been proposed as alternative explanations, but we cannot yet tell which of these models are correct, or their relative contributions.
@@
Numerous studies published prior to the release of Gaia DR3 have attempted to estimate the number of BHs that will be discovered by Gaia (Mashian & Loeb, 2017; Breivik et al., 2017; Andrews et al., 2019; Shikauchi et al., 2020; Wiktorowicz et al., 2020; Janssens et al., 2022; Shikauchi et al., 2022; Chawla et al., 2022). Most of these papers only considered the IBE formation channel, and predictions for the number of Gaia-detectable BHs have generally become more pessimistic over time (e.g., El-Badry, 2024). So far, the Gaia mission has found far fewer BHs than predicted in any of these works. Predictions published after Gaia DR3 have mainly focused on dynamical formation scenarios, which can potentially avoid the challenges associated with the IBE channel (e.g., Tanikawa et al., 2024; Di Carlo et al., 2024).
@@
Studies so far have relied on simplified detectability estimates, rather than a realistic forward-model of the Gaia astrometric binary catalogs. This approach is unlikely to yield accurate predictions because the selection function for orbital solutions published in DR3 is complex. For example, many sources that could have received well-constrained orbital solutions were instead published with acceleration solutions. Also, the cuts used in DR3 depend on quantities that are nontrivial to predict with simplified detectability metrics, such as the uncertainties on astrometric parameters. Recently, El-Badry et al. (2024) built a forward model to generate synthetic astrometric binary catalogs for DR3 and future Gaia data releases from the Gaia scanning law. In this paper, we apply this generative model to simulated populations of BH-luminous companion (BH-LC) binaries in the Milky Way to predict the number of BHs detectable in Gaia DR3, DR4, and DR5. Specifically, we investigate synthetic populations that assume either IBE (Chawla et al., 2022) or dynamical (Di Carlo et al., 2024) formation channels for Gaia BHs. By utilizing the generative model of El-Badry et al. (2024) and calibrating with the number of Gaia BHs detected in DR3, we are able to make the most realistic predictions yet for the number of Gaia BHs that will be detected in DR4 and DR5.
@@ 2.1 Isolated Binary Evolution
In their initialization, Chawla et al. (2022) assign ages and metallicities for each binary based on the simulated galaxy m12i from the FIRE-2 simulation suite (Wetzel et al., 2016; Hopkins et al., 2018), and assign 3D Galactic positions based on the ananke framework of Sanderson et al. (2020). The metallicities are constrained to fall within the valid range of . Zero-age main sequence (ZAMS) orbital parameters are drawn from observationally motivated probability distribution functions; see Chawla et al. (2022) for details. Chawla et al. (2022) use the prescription for common-envelope evolution, setting and the value of based on the default prescription in the Appendix of Claeys et al. (2014). They assume that BH natal kicks are reduced by a factor of relative to velocities drawn from a Maxwellian distribution with km s-1 (Hobbs et al., 2005), with representing the fractional fallback of the stellar envelope set by the prescription of Fryer et al. (2012).
To generate synthetic Milky Way (MW) BH-LC populations, Chawla et al. (2022) sample with replacement from a converged population of simulated binaries. To determine the number of BH-LC systems in the present-day MW, they scale up the number of BH-LC systems in the converged population by the ratio of the total mass formed in the galaxy m12i to the total mass of single and binary stars drawn to produce the simulated BH-LC population. The sampled binaries are assigned Galactocentric positions by minimizing the difference in age and metallicity with star particles in m12i; if multiple BH–LCs are assigned to the same single-star particle, they are distributed in a sphere centered around the particle with radius inversely dependent on the local density.
@@ 2.2 Binary Evolution and Cluster Dynamics
For the dynamical formation model, we adopt the population of BH-LC systems synthesized by Di Carlo et al. (2024). In their 100-Myr long simulations of 3000 open star clusters (“set A”, see Di Carlo et al., 2020), dynamics are treated by the -body code NBODY6++GPU (Wang et al., 2015), and population synthesis by MOBSE (Mapelli et al., 2017; Giacobbo et al., 2018). The star clusters (SCs) have masses ranging from – (drawn from the young SC initial mass function (IMF) of Lada & Lada 2003) and initial half-mass radii calculated using the relationship from Marks & Kroupa (2012). The SCs have three different possible metallicities .
@@
Di Carlo et al. (2024) create synthetic MW BH-LC populations by seeding both isolated binaries (IBs) and their simulated SCs into the same m12i galaxy as Chawla et al. (2022). To derive the metallicity-dependent star formation rate (SFR), they bin the star particles in metallicity and age. Then, they draw clusters with replacement in each bin until 10% of the star formation in m12i is account for. The remaining 90% is turned into massive IBs through direct conversion into star particles in each bin; for details, we direct the reader to Di Carlo et al. (2024). To match the original cluster simulations, the IBs are given absolute metallicities .
@@
For the purpose of comparing formation channels, we only consider the population of BH-LC systems dynamically formed in the SCs of Di Carlo et al. (2024) when using the forward model of El-Badry et al. (2024) to make predictions for detectable Gaia BHs. Nevertheless, to gain insight into the effect of differing input assumptions, we compare the IBE population of Di Carlo et al. (2024) against that of Chawla et al. (2022) in Section 4.3. We note that removing systems with ages Myr from the IBE population of Chawla et al. (2022) reduces its size by , but has a larger effect on the detectable population. We discuss the effects of this cut further in Section 3.2.
@@ 3.7 Searches for BHs with Acceleration Solutions
where is the uncertainty-normalized parallax, is the solution significance, and the statistic represents the goodness-of-fit.
@@ 4.3 Importance of the Adopted Binary Evolution Model
We plot orbital period versus distance and eccentricity for the dynamically assembled BH-LC systems of Di Carlo et al. (2024) in Figure 13. In the top row, we plot primordial binaries in orange and dynamically exchanged binaries in blue. We find that most of the detected Gaia BHs are formed via dynamical exchanges. In the bottom row, we plot binaries that escaped from their open clusters in gold and binaries retained in their open clusters in purple. We find that most detected Gaia BHs are retained for at least 100 Myr in the open clusters in which they formed. However, since many systems will likely still escape at later times, this does not necessarily indicate that (if BH-LC systems preferentially form via the dynamical channel) we should expect to discover many Gaia BHs in open clusters in future Gaia data releases.
In all panels, there is a steep drop off in the period distribution at days, which reflects the fact that the input period distribution from Sana et al. (2012) only extends to days. However, Figure 13 shows that most of the binaries are exchanged, not primordial. This may appear counter-intuitive, since one may expect exchanges to erase memory of the input period distribution. However, the models from Di Carlo et al. (2024) predict that the post-exchange period is usually similar to or shorter than the pre-exchange period (see also Sigurdsson & Phinney, 1993), so the pre-exchange period distribution, and thus the assumed binary evolution, is actually quite important.

### v2

@@ Abstract
Astrometry from Gaia has enabled discovery of three dormant black holes (BHs) in au-scale binaries. Numerous models have been proposed to explain their formation, including several that have forecasted Gaia detections. However, previous works have used simplified detectability metrics that do not capture key elements of the Gaia astrometric orbit selection function. We apply a realistic forward-model of Gaia astrometric orbit catalogs to BH binary populations generated through (a) isolated binary evolution (IBE) and (b) dynamical formation in star clusters. For both formation channels, we analyze binary populations in a simulated Milky Way-like galaxy with a realistic metallicity-dependent star formation history and 3D dust map. We generate epoch astrometry for each binary from the Gaia scanning law and fit it with the cascade of astrometric models used in Gaia DR3. The IBE model of Chawla et al. (2022) predicts that no BH binaries should have been detected in DR3 and thus significantly underpredicts the formation rate of Gaia BHs. In contrast, the dynamical model of Di Carlo et al. (2024) overpredicts the number of BHs receiving DR3 orbital solutions by a factor of . The two models predict very different orbital period distributions, with the IBE model predicting only binaries that avoided common envelope evolution and have d to be detectable, and the dynamical formation model predicting a period distribution that is roughly log-uniform. Adopting the dynamical channel as a fiducial model and rescaling by a factor of to match DR3, we predict that BH binaries will be detected in Gaia DR4, representing of Milky Way BHs with luminous companions in au-scale orbits.
@@ 1 Introduction
The formation of Gaia BH binaries remains uncertain. Isolated binary evolution (IBE) models struggle to reproduce the systems’ wide orbits. Their current orbital separations are significantly smaller than the predicted maximum radii of their BH progenitors, suggesting that their luminous stars would have interacted with the progenitors when they were red supergiants. However, these binaries would also have to survive an episode of common-envelope evolution with a large donor-to-accretor mass ratio; even if the envelopes of the massive stars were successfully ejected, the final orbits are predicted to be far tighter than observed. Modified prescriptions of IBE (e.g., Kruckow et al., 2024; Gilkis & Mazeh, 2024; Kotko et al., 2024; Iorio et al., 2024), formation in a hierarchical triple (e.g., Nagarajan et al., 2024; Generozov & Perets, 2024; Li et al., 2024; Tanikawa et al., 2024b) and dynamical assembly in a star cluster (e.g., Rastello et al., 2023; Tanikawa et al., 2024a; Di Carlo et al., 2024; Fantoccoli et al., 2025; Marín Pina et al., 2024) have all been proposed as alternative explanations, but we cannot yet tell which of these models are correct, or their relative contributions.
@@
Numerous studies published prior to the release of Gaia DR3 have attempted to estimate the number of BHs that will be discovered by Gaia (e.g., Mashian & Loeb, 2017; Breivik et al., 2017; Yamaguchi et al., 2018; Yalinewich et al., 2018; Shao & Li, 2019; Andrews et al., 2019; Shikauchi et al., 2020; Wiktorowicz et al., 2020; Chawla et al., 2022; Janssens et al., 2022; Shikauchi et al., 2022; Wang et al., 2022). Most of these papers only considered the IBE formation channel, and predictions for the number of Gaia-detectable BHs have generally become more pessimistic over time (e.g., El-Badry, 2024). So far, the Gaia mission has found far fewer BHs than predicted in any of these works. Predictions published after Gaia DR3 have mainly focused on dynamical formation scenarios, which can potentially avoid the challenges associated with the IBE channel (e.g., Tanikawa et al., 2024a; Di Carlo et al., 2024).
@@
Studies so far have relied on simplified detectability estimates, rather than a realistic forward-model of the Gaia astrometric binary catalogs. This approach is unlikely to yield accurate predictions because the selection function for orbital solutions published in DR3 is complex. For example, many sources that could have received well-constrained orbital solutions were instead published with acceleration solutions (Pourbaix et al., 2022; El-Badry et al., 2024). Also, the cuts used in DR3 depend on quantities that are nontrivial to predict with simplified detectability metrics, such as the uncertainties on astrometric parameters (Halbwachs et al., 2023). Recently, El-Badry et al. (2024) built a forward model to generate synthetic astrometric binary catalogs for DR3 and future Gaia data releases from the Gaia scanning law. In this paper, we apply this generative model to simulated populations of BH-luminous companion (BH-LC) binaries in the Milky Way to predict the number of BHs detectable in Gaia DR3, DR4, and DR5. Specifically, we investigate synthetic populations that assume either IBE (Chawla et al., 2022) or dynamical (Di Carlo et al., 2024) formation channels for Gaia BHs. By utilizing the generative model of El-Badry et al. (2024) and calibrating with the number of Gaia BHs detected in DR3, we are able to make the most realistic predictions yet for the number of Gaia BHs that will be detected in DR4 and DR5.
@@ 2.1 Isolated Binary Evolution
In their initialization, Chawla et al. (2022) assign ages and metallicities for each binary based on the simulated galaxy m12i from the FIRE-2 simulation suite (Wetzel et al., 2016; Hopkins et al., 2018), and assign 3D Galactic positions based on the ananke framework of Sanderson et al. (2020), which generates mock catalogs of synthetic stars from cosmological baryonic simulations. The metallicities are constrained to fall within the valid range of . Zero-age main sequence (ZAMS) orbital parameters are drawn from observationally motivated probability distribution functions; see Chawla et al. (2022) for details. Chawla et al. (2022) use the prescription for common-envelope evolution, setting and the value of based on the default prescription in the Appendix of Claeys et al. (2014). They assume that BH natal kicks are reduced by a factor of relative to velocities drawn from a Maxwellian distribution with km s-1 (Hobbs et al., 2005), with representing the fractional fallback of the stellar envelope set by the prescription of Fryer et al. (2012).
To generate synthetic Milky Way (MW) BH-LC populations, Chawla et al. (2022) sample with replacement from a converged population of simulated binaries, where convergence refers to an iterative process in which the shapes of binary parameter distributions stabilize as the simulated population grows (Breivik et al., 2020). To determine the number of BH-LC systems in the present-day MW, they scale up the number of BH-LC systems in the converged population by the ratio of the total mass formed in the galaxy m12i to the total mass of single and binary stars drawn to produce the simulated BH-LC population. The sampled binaries are assigned Galactocentric positions by minimizing the difference in age and metallicity with star particles in m12i; if multiple BH–LCs are assigned to the same single-star particle, they are distributed in a sphere centered around the particle with radius inversely dependent on the local density.
@@ 2.2 Binary Evolution and Cluster Dynamics
For the dynamical formation model, we adopt the population of BH-LC systems synthesized by Di Carlo et al. (2024). In their 100-Myr long simulations of 3000 open star clusters (“set A”, see Di Carlo et al., 2020), dynamics are treated by the -body code NBODY6++GPU (Wang et al., 2015), and population synthesis by MOBSE (Mapelli et al., 2017; Giacobbo et al., 2018). The star clusters (SCs) have masses ranging from – (drawn from the young SC initial mass function (IMF) of Lada & Lada 2003) and initial half-mass radii calculated using the relationship from Marks & Kroupa (2012). The SCs have three different metallicities and .
@@
Di Carlo et al. (2024) create synthetic MW BH-LC populations by seeding both isolated binaries (IBs) and their simulated SCs into the same m12i galaxy as Chawla et al. (2022). To derive the metallicity-dependent star formation rate (SFR), they bin the star particles in metallicity and age. Then, they draw clusters with replacement in each bin until 10% of the star formation in m12i is account for. The remaining 90% is turned into massive IBs through direct conversion into star particles in each bin; for details, we direct the reader to Di Carlo et al. (2024). To match the original cluster simulations, the IBs are given absolute metallicities (i.e., representing the mass fraction of metals).
@@
For the purpose of comparing formation channels, we only consider the population of BH-LC systems formed in the SCs of Di Carlo et al. (2024) when using the forward model of El-Badry et al. (2024) to make predictions for detectable Gaia BHs. This includes both primordial binaries and dynamically exchanged binaries in those SCs (see Section 4.3). Nevertheless, to gain insight into the effect of differing input assumptions, we compare the IBE population of Di Carlo et al. (2024) against that of Chawla et al. (2022) in Section 4.3. We note that removing systems with ages Myr from the IBE population of Chawla et al. (2022) reduces its size by , but has a larger effect on the detectable population. We discuss the effects of this cut further in Section 3.2.
@@ 3.7 Searches for BHs with Acceleration Solutions
where is the uncertainty-normalized parallax, is the solution significance, and the statistic represents the goodness-of-fit (El-Badry et al., 2024).
@@ 4.3 Importance of the Adopted Binary Evolution Model
We plot orbital period versus distance and eccentricity for the BH-LC systems formed in the SCs of Di Carlo et al. (2024) in Figure 13. In the top row, we plot primordial binaries in orange and dynamically exchanged binaries in blue. We find that most Gaia BHs are formed via dynamical exchanges. In the bottom row, we plot binaries that escaped from their open clusters in gold and binaries retained in their open clusters in purple. We find that most Gaia BHs are retained for at least 100 Myr in the open clusters in which they formed. However, since many systems will likely still escape at later times, this does not necessarily indicate that (if BH-LC systems preferentially form via the dynamical channel) we should expect to discover many Gaia BHs in open clusters in future Gaia data releases. Moreover, even if Gaia BHs are retained in open clusters, they could be difficult to detect due to the effects of crowding.
In all panels, there is a steep drop off in the period distribution at days, which reflects the fact that the input period distribution from Sana et al. (2012) only extends to days. However, Figure 13 shows that most of the binaries are exchanged, not primordial. This may appear counter-intuitive, since one may expect exchanges to substantially alter the input period distribution. However, the models from Di Carlo et al. (2024) predict that the post-exchange period is usually similar to or shorter than the pre-exchange period (see also Sigurdsson & Phinney, 1993), so the pre-exchange period distribution, and thus the assumed binary evolution, is actually quite important.